OpenStudy (anonymous):

evaluate the given integral e^5x

OpenStudy (kirbykirby):

\[\int e^{5x} dx\] Let \(u=5x\), then \(du=5dx\) => \(\frac{du}{5}=dx\) Now, you get \[\int e^u \frac{1}{5}du\] \[=\frac{1}{5}\int e^u du\] \[=\frac{1}{5}e^u+C\] Substitute back your substitution "u=5x": \(\frac{1}{5}e^{5x}+C\)